Professional Headhunters, Inc. (PHI), a job placement company, operates in the northeastern United States. During Year 1, the company earned $145,000 in revenue by providing services to customers. However, it collected only $120,000 of the revenue in cash. PHI expected to collect the remaining $25,000 in Year 2. In addition, PHI incurred $80,000 of expenses. However, by the end of Year 1, PHI had paid only $75,000 of the cash owed for expenses because it had not yet paid $5,000 to employees who had worked during Year 1 but had not been paid by the end of the year. PHI expected to pay the $5,000 in cash to the employees during Year 2. Based on this information alone, determine the amount of net income, total assets, and total liabilities PHI should report on its Year 1 financial statements.
Answer #1

Net income = Service revenue - Expenses

= 145,000 - 80,000

= $65,000

Cash balance = Cash collected from customers - Cash paid for expenses

= 120,000 - 75,000

= $45,000

Total assets = Cash + Accounts receivables

= 45,000 + 25,000

= $70,000

Total liabilities = Salaries payable

= $5,000